OpenStudy (anonymous):

16! / 4! 12! =??? A. .0238 B. 1,820 C. 2,380

OpenStudy (anonymous):

you can cancel out 12! from the top and bottom and you get 16*15*14*13/4*3*2*1, then cancel out some more to get 4*5*7*13=140*13=1820

OpenStudy (lgbasallote):

for better visualization :) \[\huge \frac{16!}{4!12!} \implies \frac{16 \times 15 \times 14 \times 13 \times 12!}{4!12!}\]

OpenStudy (anonymous):

thank you guys! :)